OpenStudy (anonymous):

What is the quotient of five over eighteen divided by five over twelve in simplest form

OpenStudy (mokeira):

\[\frac{ 5 }{ 18 }\div \frac{ 5 }{ 12 }\] \[\frac{ 5 }{ 18 }\times \frac{ 12 }{ 5 }\] \[\frac{ 60 }{ 90 }\] \[\frac{ 2 }{ 3 }\]